OpenStudy (kainui):

I'm just never satisfied by any "proofs" of the pythagorean theorem. Can someone give me insight why the sum of areas of similar shapes on two sides equals the area of another similar shape on the diagonal side? Seriously, why? It's just completely no intuition there.

OpenStudy (kainui):

Those are exactly what _don't_ help.

OpenStudy (kainui):

Nothing about rearranging cut out squares is intuitive. I'm looking for a deeper answer. If for some reason the pythagorean theorem was a^3+b^3=c^3 and they drew it all out with volumes and it matched up I'd be equally as dissatisfied.

OpenStudy (anonymous):

to tell you the truth much math in that era was drawn and then solve theoretically so i dont think you can really get away from drawing squares to = hypot square

OpenStudy (kainui):

Yeah, I don't think either of you get where I'm coming from. I've seen many different things like this, but none of them make it intuitive. For instance, I like this:  \[(a+b)^2=4*\frac{ 1 }{ 2 }ab+c^2\] it follows, a^2+b^2=c^2 But that doesn't EXPLAIN anything, only SHOWS that it is true. WHY?

OpenStudy (primeralph):

@Kainui Hold on.

OpenStudy (primeralph):